OpenStudy (anonymous):

FIND THE RECIPROCAL: 1) 0.66 2) -2.23

OpenStudy (anonymous):

1. 0.66=2/3 3/2.

OpenStudy (anonymous):

thank you

OpenStudy (stormfire1):

1) .66 = .66/1 so the reciprocal is 1/.66 2) -2.23 = -2.23/1 so the reciprocal is -1/2.23

OpenStudy (anonymous):

ohhh. i see now. thanks (:

OpenStudy (anonymous):

\[\text{1) }\left(\frac{66}{100}\right)^{-1}=\frac{50}{33} \]\[2)\text{ }\left(-\frac{223}{100}\right)^{-1}=-\frac{100}{223}\]